OpenStudy (m_m13):

will medal and fan What is the point-slope form of the equation of a line that passes through the point (3, 5) and has a slope of 2?

OpenStudy (m_m13):

A. y + 3 = 2(x + 5) B. y - 3 = 2 (x - 5) C. y - 5 = 2 (x - 3) D. y - 5 = -2 (x - 3)

OpenStudy (mathmate):

Hint: Point slope form of a line with slope m passing through (x0,y0) is y-y0=m(x-x0).

OpenStudy (m_m13):

@mathmate would it be C then?

OpenStudy (mathmate):

Agree. In the future, please copy and paste the actual answer instead of the letter to avoid misunderstanding, and that also helps you confirm your own answer.

OpenStudy (m_m13):

okay thank you